OpenStudy (anonymous):

Ryan made 104/7 gallons of cider for his party. His guests drank 61/2 gallons in all. About how much cider does Ryan have left?

OpenStudy (mathstudent55):

Are the numbers \( \dfrac{104}{7} \) and \(\dfrac{61}{2} \), which is what you wrote, or are the numbers \( 10 \dfrac{4}{7}\) and \(6 \dfrac{1}{2} \) which maybe is what you meant to write?

OpenStudy (anonymous):

10 4/7 and 6 1/2

OpenStudy (mathstudent55):

Ok. You need to subtract two mixed numerals. 1) One way of doing it is to change each mixed numeral into a fraction. Then you need to rewrite the two fractions with a common denominator. Then you subtract the fractions. Then you convert the result to a mixed numeral. 2) The other way of doing it is to leave the numbers as mixed numerals. Subtract the fraction parts using a common denominator and borrowing if necessary. Subtract the number parts.

OpenStudy (mathstudent55):

Are you familiar with either method?